# DepthVanish: Optimizing Adversarial Interval Structures for Stereo-Depth-Invisible Patches

### v1

@@ 5 Methodology
In this work, we build upon our novel findings in Sec. 3 and propose realizing the attack goal in Eq. (4) by exploiting the attack capability of interval spacing. However, this is a non-trivial problem since ❶ Eq. (4) requires the patch’s ground-truth depth to be close but Fig. 2 (a) indicates interval spacing performs weak when the patch is deployed closely. Moreover, ❷ the robustness to rotation is a critical requirement for the patch to be physically attack effective. Yet we observed in Fig. 3 that the robustness provided by the naive interval strategy is rather limited especially for Y axis. It is obvious that an advanced interval spacing strategy is required to realize our attack goal in Eq. (4).
Fundamentally, interval spacing induces a mask $\mathbf{M}$ that partitions the patch $\mathbf{P}$ into interval structure $\mathbf{P}_{s}=\mathbf{M}\odot\mathbf{P}$ and texture structure $\mathbf{P}_{t}=(1-\mathbf{M})\odot\mathbf{P}$, such that $\mathbf{P}=\mathbf{P}_{s}+\mathbf{P}_{t}$. Hence, we propose to optimize these components to reveal their adversarial effects. Since the interval spacing—and thus the mask $\mathbf{M}$—is predefined (Sec. 3), we focus on optimizing the texture structure $\mathbf{P}_{t}$, which consists of tiled texture elements $\mathbf{E}$, forming the basis of our Grid-based Attack. We then introduce the Depth Vanish Attack, which jointly optimizes both $\mathbf{P}_{s}$ and $\mathbf{P}_{t}$ for maximal effect.
@@ 5.1 Grid-based Attack
In general, it is straightforward to setup an optimization pipeline for optimizing the texture element with grid intervals, where the patch is formed by repeating the texture elements over the grid. Fundamentally, there two main aspects need to be considered: ❶ the physical constraint required for the texture element to form a patch and ❷ the objective function adopted for optimization.
Given our primary goal is to achieve physical attack effectiveness, the patch must comply with the physical geometry constraints during the optimization. In specific, given a user pre-defined patch physical size $(u,v)$ and physical distance to the camera $e$ in meters, we first find the corresponding pixel size of the patch $(h_{p},w_{p})$ with the help of stereo calibration information (See details in Sec. 6.1). Then, we empirically apply the optimal interval width $o$ and repetition times $k$ from Sec. 3 to determine the texture element size $(h_{t},w_{t})$ for $\mathbf{E}$ as,
$h_{t}=\frac{h_{p}-k\cdot o}{k+1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ w_{t}=\frac{w_{p}-k\cdot o}{k+1}.$ (5)
Based on the size of the texture element, the texture component $\mathbf{P}_{t}$—and consequently the full patch $\mathbf{P}$—is constructed by tiling the base texture unit $\mathbf{E}$ in a regular grid pattern. With the correctly assembled and deployed grid-based patch, we set the optimizing objective function as regional mean square error (rMSE) which is formulated as
@@
Let $\mathcal{R}$ be the set of $k\times k$ grid locations on where $\mathbf{E}$ is repeated. The texture element is updated with average gradients: $\mathbf{E}\leftarrow\mathbf{E}-\eta\cdot\frac{1}{|\mathcal{R}|}\sum_{(i,j)\in\mathcal{R}}\nabla_{\mathbf{E}}\mathcal{L}_{rMSE}^{(i,j)},$ where $\eta$ is the learning rate. Gradients are only applied to the repeated texture regions; the interval areas remain untouched.
@@ 5.2 Depth Vanish Attack
As we will see in Fig. 5, the above grid-based optimization can successfully realize attack against various stereo systems but the results are still far from our attack goal defined in Eq. (4). Thus we further consider to optimize the interval structure $\mathbf{P}_{s}$ at the same time as updating the texture element $\mathbf{E}$. Practically, optimizing the interval structure on patch level will break the texture repetitions as the interval will be updated to have irregular size. To keep the repetitions and incorporate the interval’s attack capability, we propose to optimize the interval structure within the texture element, i.e., element level, then following berger2022stereoscopic repeat the optimized texture element $\mathbf{E}$ forming the final patch.
Same to grid-based attack, given a user pre-defined patch physical size $(u,v)$ and physical distance to the camera $e$ in meters, we first find the corresponding pixel size of the patch $(h_{p},w_{p})$. Then we calculate the texture element size $(h_{t},w_{t})$ by simply dividing $(h_{p},w_{p})$ to the repetition times $k$. In order to optimize the texture element so that the interval structure integrated as part of the texture, we propose to regularize the texture element $\mathbf{E}$ during optimization with two objectives. First, we directly cast entropy constraint on the texture element for regularizing its values to be binary, so that a crisp separation is formed to serve as the required interval structure.
$\mathcal{L}_{entropy}=\frac{1}{h_{t}\cdot w_{t}}\sum_{i=1}^{h_{t}}\sum_{j=1}^{w_{t}}-\mathbf{E}_{ij}log(\mathbf{E}_{ij}+\epsilon)-(1-\mathbf{E}_{ij})log(1-\mathbf{E}_{ij}+\epsilon),$ (7)
However, we experimentally found that the texture element cannot form a clear pattern with only entropy regularization. As a result, we further integrate the total variation loss to penalizes local pixel-level variation, encouraging the formation of smooth areas.
@@
$\mathcal{L}=\mathcal{L}_{rMSE}+\alpha*\mathcal{L}_{entropy}+\beta*\mathcal{L}_{tv}$ (9)
where $\alpha$ and $\beta$ are the hyper-parameters balancing the sharp border and coherent region requirements. Hence, we update with, $\mathbf{E}\leftarrow\mathbf{E}-\eta\cdot\frac{1}{|\mathcal{R}|}\sum_{(i,j)\in\mathcal{R}}\nabla_{\mathbf{E}}\mathcal{L}^{(i,j)},$ where $\eta$ is the learning rate.
@@
During the optimization of grid-based and depth vanish attack, we keep adopt a patch with a physical size $(u,v)=(0.891m,1.26m)$ and specify the physical ground-truth depth $e=5m$. To assemble the texture element into patch, we empirically set the repetition as 5 for horizontal and 4 for vertical, i.e., $k=(4,5)$. For the optimization and corresponding evaluation results with different patch physical setup, we provide them in the supplemental material. When the patch is optimize as grid-based attack, the optimal interval size $o=10$px from Sec. 3 is applied. As for the loss weights adopted during the depth vanish attack, we keep setting $\alpha=0.1$ and $\beta=10$. Please find more details of implementation for both grid-based and depth vanish attack in the supplement.
@@ 6.1 Experimental Setup
For the evaluation of digital attack effectiveness, we adopt the stereo images from KITTI scene flow (KITTI-scene) kittisceneflow and DrivingStereo yang2019drivingstereo datasets. Both datasets composed of stereo images for urban traffic scenes where the image size of KITTI-scene is (1242,375) and DrivingStereo is (1758, 800). In more detail, we adopt the four sub-sets of DrivingStereo that captured under different weather conditions (i.e., sunny, foggy, rainy, cloudy) where we report the attack performance for each of them respectively. Following berger2022stereoscopic , 40 stereo image pairs for each (sub-)dataset are selected to verify the effectiveness of different patches. For the physical evaluation, we manually capture stereo images with i3DStreoid  [^1] http://stereo.jpn.org/eng/iphone/help/index.html where various safety critical situations are considered. We refer readers to the supplemental material for the details of how the physical stereo images are captured and the pipeline we adopted for physical deployment.
@@
Following wong2021stereopagnosia ; berger2022stereoscopic , we apply our attack method to PSMNet ChangC18PSMN , DeepPruner duggal2019deeppruner and AANet xu2020aanet for validating the general attack effectiveness. Moreover, we empirically found that they are out-of-date and easily to be disturbed, thus we further select RAFT-Stereo lipson2021raft and STereo TRansformer (STTR) li2021revisiting which represents the most promising iterative optimization-based methods and transformer-base methods as our main attack targets. For the detail hyper-parameter setting and the pretrained checkpoint adopted during the attack, please find all of them in supplemental material.
@@
During the digital optimization and evaluation, the patch needs to be placed inside the scene following the physical constraints. To achieve this, we apply the calibration information provided by the KITTI and DrivingStereo dataset. In specific, given a patch with predefined physical size in meters, we first set the homogeneous 3D position of the patch’s corners regarding to the reference camera coordinate system. Then we calculate the corresponding pixel coordinates with the help of the rectified projection and rotation matrix. The full calculation is detailed in supplement.
@@
Following the convention, we adopt bad pixel error (D1-error) and End-Point Error (EPE) for evaluating the prediction performance and they are specifically calculated as
[…]
where the bad pixel is one that satisfy $|\textbf{d}_{pred}-\textbf{d}_{gt}|>\text{max}(3,0.05\cdot\textbf{d}_{gt})$. To evaluate patch attack effectiveness, we first follow Eq. (4) to set the ground-truth disparity of the patch as $\textbf{d}_{gt}=\textbf{c}$. Then, we define the bad pixels as those satisfying $|\textbf{d}_{pred}-\textbf{c}|>\text{max}(3,0.05\cdot\textbf{d}_{gt})$ and $|\textbf{d}_{pred}-\textbf{0}|<\frac{\textbf{c}}{k}$, where $k$ defines how many times deeper than the actual depth will a patch be considered to be attack effective. In summary, we report the average D1-error and EPE with standard deviation where higher values indicate better attack performance.
@@ 6.2 Digitally Attack Stereo Estimator
Setting: Due to the lack of existing works on attacking stereo matching with patch, we apply the results from existing digital attack studies (i.e., Stereoscopic berger2022stereoscopic and Stereopagnosia wong2021stereopagnosia ) as patches and deploy them into the scene as the first set of baselines. However, it should be noted that such comparison is not fair enough as existing works berger2022stereoscopic ; wong2021stereopagnosia are not specifically designed for patch attack. Thus we further setup our own baseline (i.e., grid-based patch from Sec. 5.1) for a fair comparison.
Results: ❶ We report the attack results for the five attack target models on KITTI-scene dataset in Tab. 1. It can be seen that existing digital attacks are ineffective under the patch attack setup, while our Grid-based Patch significantly outperforms them. Notably, DepthVanish achieves strong attack performance, especially against DeepPruner, RAFT-Stereo and STTR. We illustrate the results on DrivingStereo dataset in Fig. 4. It is obvious that similar attack performance can be observed on all the four sub-sets. ❷ In addition to the standard evaluation, Fig. 5 shows a KITTI sample comparison. Compared to the Clean Depth, we first note that existing attack works fail to mislead all the five target models, where only the Stereoscopic Patch shows slightly influence on PSMNet. However, as the results shown in the last column our DepthVanish patch can cast strong influence where the patch almost disappear. More surprisingly, our patch enjoys significant transferability over models where the patch optimized with PSMNet shows strong attack effect on other four models. From our experimental experience, all the patches with such clear interval patterns are model transferable for which we credit such capability to the insights we analyzed in Sec. 3. Please refer to the supplemental material for the comprehensive experimental results of attack transferability.
@@ 6.3 Physically Attack Stereo Estimator
In this section, we conduct physical evaluation for our DepthVanish patches that optimized with different stereo estimators to highlight the importance and emergency of research on stereo matching reliability. As shown in Fig. 6, we host our DepthVanish patches on a white board for the purpose of contrasting the depth disturbance. ❶ It can be observed from the results that our DepthVanish patch consistently preserves its attack effectiveness after deployed into the physical environment. Compared to the Clean Depth, the board region occupied by our DepthVanish patches are predicted as far away in general. ❷ However, it should be noted that the induced depth error are limited compared to the digital effectiveness in Fig. 5. We ascribe such performance degradation to the lighting variation and imprecise photo capturing process, where the left and right images are captured manually in separate. Therefore, we further conduct evaluation for our patch against commercial stereo depth camera in the next. In summary, despite of the imprecise stereo image capturing process, our DepthVanish attack still keeps successfully attacking SOTA DNN-based stereo depth estimation models with consistency.
@@ 6.4 Attack Commercial Stereo System
To further assess the practicality and robustness of our DepthVanish patch, we evaluate its performance on a commercial stereo camera system, specifically the Intel RealSense D435i depth camera. We focus on evaluating the patch’s robustness from three aspects: model generalization, geometric transformation, and distance variation. ❶ Model generalization: we deploy the patch, optimized on the KITTI dataset, against five stereo models without any extra adaptation. As shown in Fig. 7, the patch consistently disrupts depth predictions across all models, demonstrating strong cross-model transferability. ❷ Geometric robustness: we physically rotate the patch along the X and Y axes. As visualized in Fig. 8, the patch remains effective under different viewing angles, confirming its robustness to rotation. ❸ Distance robustness: our method also shows robustness under varying distances. Corresponding visual results are provided in the supplementary material.

### v2

@@ 5 Methodology
In this work, we build upon our novel findings in Sec. 3 and propose realizing the attack goal in Eq. (4) by exploiting the attack capability of interval spacing. However, this is a non-trivial problem since ❶ Eq. (4) requires the patch’s ground-truth depth to be close but Fig. 2 (a) indicates that interval spacing exerts only a limited adversarial effect when the patch is deployed closely. Moreover, ❷ the robustness against rotation is a critical requirement for the patch to be physically attack effective. Yet we observed in Fig. 3 that the robustness provided by the naive interval strategy is rather limited especially against the rotation of Y axis. As a result, it is obvious that an advanced interval spacing strategy is required to realize our attack goal as defined in Eq. (4).
Fundamentally, interval spacing induces a mask $\mathbf{M}$ that partitions the patch $\mathbf{P}$ into interval structure $\mathbf{P}_{s}=\mathbf{M}\odot\mathbf{P}$ and texture content $\mathbf{P}_{t}=(1-\mathbf{M})\odot\mathbf{P}$, such that $\mathbf{P}=\mathbf{P}_{s}+\mathbf{P}_{t}$. Hence, we propose to optimize these components to reveal their adversarial effects. Beginning with the naive interval spacing strategy, and thus the mask $\mathbf{M}$, in Sec. 3, we first focus on optimizing the texture content $\mathbf{P}_{t}$, which composed of tiled texture elements $\mathbf{E}$, forming the basis of our Grid-based Attack. We then introduce the DepthVanish Attack, which jointly optimizes both $\mathbf{P}_{s}$ and $\mathbf{P}_{t}$ for maximal effect.
@@ 5.1 Grid-based Attack
In general, it is straightforward to setup an optimization pipeline for optimizing the texture element with grid intervals, where the patch is formed by repeating the texture elements over the grid. Fundamentally, there two main aspects that need to be considered: ❶ the physical constraint required for the texture element to form a patch and ❷ the objective function adopted for optimization.
Given our primary goal is to achieve physical attack effectiveness, the patch must comply with the physical geometry constraints during the optimization. Specifically, given a user pre-defined physical patch size $(u,v)$ and physical distance to the camera $e$ in meters, we first find the corresponding pixel size of the patch $(h_{p},w_{p})$ with the help of stereo calibration information (See details in Sec. 6.1). Then, we empirically adopt the optimal interval width $o$ and number of repetition $k$ from Sec. 3 to determine the texture element $\mathbf{E}$ size $(h_{t},w_{t})$ as
$h_{t}=\frac{h_{p}-k\cdot o}{k+1},\penalty 10000\ \penalty 10000\ \penalty 10000\ w_{t}=\frac{w_{p}-k\cdot o}{k+1}.$ (5)
Based on the size of the texture element, the texture component $\mathbf{P}_{t}$, and consequently the full patch $\mathbf{P}$, is constructed by tiling the base texture unit $\mathbf{E}$ in a regular grid pattern as illustrated in Fig. 2(b) (bottom right). With the correctly assembled and deployed grid-based patch, we set the optimizing objective function as regional mean square error (rMSE) which is formulated as
@@
Let $\mathcal{R}$ be the set of $k\times k$ grid locations on where $\mathbf{E}$ is repeated. The texture element is updated with average gradients: $\mathbf{E}\leftarrow\mathbf{E}-\eta\cdot\frac{1}{|\mathcal{R}|}\sum_{(i,j)\in\mathcal{R}}\nabla_{\mathbf{E}}\mathcal{L}_{rMSE}^{(i,j)},$ where $\eta$ is the learning rate. Gradients are only applied to the repeated texture regions while the interval areas remain untouched.
@@ 5.2 DepthVanish Attack
As we will see in Fig. 5, the above grid-based optimization can successfully mount an attack against various stereo systems but the results are still far from our attack goal defined in Eq. (4). Thus we further consider optimizing the interval structure $\mathbf{P}_{s}$ simultaneously during the updating of the texture element $\mathbf{E}$. Practically, optimizing the interval structure on patch level will break the texture repetitions as the interval will be updated to have irregular size. To keep the repetitions and incorporate the interval’s attack capability, we propose to jointly optimize the interval structure within the texture element and, following berger2022stereoscopic , tile the optimized texture element $\mathbf{E}$ to form the final patch.
Same to grid-based attack, given a user pre-defined patch physical size $(u,v)$ and physical distance to the camera $e$ in meters, we first find the corresponding pixel size of the patch $(h_{p},w_{p})$. Then we calculate the texture element size $(h_{t},w_{t})$ by simply dividing $(h_{p},w_{p})$ to the repetition times $k$. In order to optimize the texture element so that the interval structure integrated as part of the texture, we propose to regularize the texture element $\mathbf{E}$ during optimization with two objectives. First, we directly cast entropy constraint on the texture element for regularizing its values to be binary, so that a crisp separation is formed to serve as the required interval structure:
$\mathcal{L}_{entropy}=\frac{1}{h_{t}\cdot w_{t}}\sum_{i=1}^{h_{t}}\sum_{j=1}^{w_{t}}-\mathbf{E}_{ij}log(\mathbf{E}_{ij}+\epsilon)-(1-\mathbf{E}_{ij})log(1-\mathbf{E}_{ij}+\epsilon).$ (7)
However, we experimentally found that the texture element cannot form a clear pattern with only entropy regularization. As a result, we further integrate the total variation loss to penalizes local pixel-level variation, encouraging the formation of smooth areas:
@@
$\mathcal{L}=\mathcal{L}_{rMSE}+\alpha*\mathcal{L}_{entropy}+\beta*\mathcal{L}_{tv},$ (9)
where $\alpha$ and $\beta$ are the hyper-parameters balancing the sharp border and coherent region requirements. Hence, we update with $\mathbf{E}\leftarrow\mathbf{E}-\eta\cdot\frac{1}{|\mathcal{R}|}\sum_{(i,j)\in\mathcal{R}}\nabla_{\mathbf{E}}\mathcal{L}^{(i,j)}$ where $\eta$ is the learning rate.
@@
During the optimization for the Grid-based and DepthVanish attacks, we use a patch with a physical size $(u,v)=(0.891\penalty 10000\ m,1.26\penalty 10000\ m)$ and specify the physical ground-truth depth $e=5\penalty 10000\ m$. To assemble the texture element into a patch, we empirically set the number of repetition as 5 for horizontal and 4 for vertical, i.e., $k=(4,5)$. For the optimization and corresponding evaluation results with different patch physical setup, we provide them in the supplemental material. When the patch is optimized as grid-based attack, the optimal interval size $o=10\penalty 10000\ px$ from Sec. 3 is applied. As for the loss weights adopted during the depth vanish attack, we keep setting $\alpha=0.1$ and $\beta=10$. Please find more details of implementation for both Grid-based and DepthVanish attack in the supplemental material.
@@ 6.1 Experimental Setup
For the evaluation of digital attack effectiveness, we adopt the stereo images from KITTI scene flow (KITTI-scene) kittisceneflow and DrivingStereo yang2019drivingstereo datasets. Both datasets are composed of stereo images of urban traffic scenes where the image size of KITTI-scene is (1242, 375) and DrivingStereo is (1758, 800). In more detail, we adopt the four sub-sets of DrivingStereo that were captured under different weather conditions (i.e., sunny, foggy, rainy, cloudy) where we report the attack performance for each of them respectively. Following berger2022stereoscopic , 40 stereo image pairs for each (sub-)dataset are selected to verify the effectiveness of different patches. For the physical evaluation, we manually capture stereo images with i3DStreoid  [^1] http://stereo.jpn.org/eng/iphone/help/index.html where various safety critical situations are considered. We refer readers to the supplemental material for the details of how the physical stereo images are captured and the pipeline we adopted for physical deployment.
@@
Following wong2021stereopagnosia ; berger2022stereoscopic , we apply our attack method to PSMNet ChangC18PSMN , DeepPruner duggal2019deeppruner and AANet xu2020aanet for validating the general attack effectiveness. Moreover, we empirically found that they are out-of-date and can be easily disturbed, thus we further select RAFT-Stereo lipson2021raft and STereo TRansformer (STTR) li2021revisiting which represent the promising iterative optimization-based methods and transformer-base methods as our main attack targets. For the detail hyper-parameter setting and the pretrained checkpoint adopted during the attack, please find all of them in supplemental material.
@@
During the digital optimization and evaluation, the patch needs to be placed inside the scene according to physical constraints. To achieve this, we apply the calibration information provided by the KITTI and DrivingStereo dataset. In specific, given a patch with a predefined physical size in meters, we first set the homogeneous 3D coordinates of the patch’s corners with respect to the reference camera coordinate system. Then we calculate the corresponding pixel coordinates with the help of the rectified projection and rotation matrix. The full calculation is detailed in supplement.
@@
Following the convention, we adopt bad pixel error (D1-error) and End-Point Error (EPE) for evaluating the prediction performance which are calculted as follows:
$\text{D1}=\frac{\#\penalty 10000\ \text{of bad pixels}}{\#\penalty 10000\ \text{of total pixels}}\times 100\%,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{EPE}=\frac{1}{N}\sum_{i=1}^{N}|\textbf{d}^{i}_{pred}-\textbf{d}^{i}_{gt}|,$ (10)
where the bad pixel is one that satisfy $|\textbf{d}_{pred}-\textbf{d}_{gt}|>\text{max}(3,0.05\cdot\textbf{d}_{gt})$. To evaluate patch attack effectiveness, we first follow Eq. (4) to set the ground-truth disparity of the patch as $\textbf{d}_{gt}=\textbf{c}$. Then, we define the bad pixels as those satisfying $|\textbf{d}_{pred}-\textbf{c}|>\text{max}(3,0.05\cdot\textbf{d}_{gt})$ and $|\textbf{d}_{pred}-\textbf{0}|<\frac{\textbf{c}}{n}$, where $n$ defines how many times deeper than the actual depth will a patch be considered to be attack effective. In summary, we report the average D1-error and EPE with standard deviation where higher values indicate better attack performance.
@@ 6.2 Digitally Attack Stereo Estimator
Setting: Due to the lack of existing works on attacking stereo matching using patches, we use the results from existing digital attack studies (i.e., Stereoscopic berger2022stereoscopic and Stereopagnosia wong2021stereopagnosia ) as patches and deploy them into the scene as the first set of baselines. However, it should be noted that such comparison is not fair enough as existing works berger2022stereoscopic ; wong2021stereopagnosia are not specifically designed for patch attack. Thus we further setup our own baseline (i.e., grid-based patch from Sec. 5.1) for a fair comparison.
Results: ❶ We report the attack results for the five attack target models on KITTI-scene dataset in Tab. 1. It can be seen that existing digital attacks are ineffective under the patch attack setup, while our Grid-based Patch significantly outperforms them. Notably, DepthVanish achieves strong attack performance, especially against DeepPruner, RAFT-Stereo and STTR. We illustrate the results on DrivingStereo dataset in Fig. 4. It is evident that similar attack performance can be observed on all four sub-sets. ❷ In addition to the standard evaluation, Fig. 5 shows a KITTI sample comparison. Compared to the Clean Depth, we first note that existing attack works fail to mislead all the five target models, where only the Stereoscopic Patch shows limited influence against PSMNet. However, as the results shown in the last column, our DepthVanish patch casts strong influence where it almost disappeared within the depth results. More surprisingly, our patch enjoys significant transferability over models where the patch optimized with PSMNet shows strong attack effect on other four models. Based our experimental experience, all patches with such clear interval patterns are transferable across models, a capability we attribute to the insights analyzed in Sec. 3. Please refer to the supplement for the comprehensive experimental results of attack transferability.
@@ 6.3 Physically Attack Stereo Estimator
In this section, we conduct physical evaluation for our DepthVanish patches that optimized with different stereo estimators to highlight the importance and emergency of research on stereo matching reliability. As shown in Fig. 6, we host our DepthVanish patches on a white board for the purpose of highlighting the depth inconsistency. ❶ It can be observed from the results that our DepthVanish patch consistently preserves its attack effectiveness after deployed into the physical environment. Compared to the Clean Depth, the board region occupied by our DepthVanish patches are predicted as far away in general. ❷ However, it should be noted that the induced depth error are limited compared to the digital effectiveness in Fig. 5. We ascribe such performance degradation to the lighting variation and imprecise photo-capturing process, where the left and right images are captured manually and separately. Therefore, we further conduct evaluation for our patch against a commercial stereo depth camera in the next section. In summary, despite of the imprecise stereo image capturing process, our DepthVanish patches successfully attack advanced DNN-based stereo estimators with consistency.
@@ 6.4 Attack Commercial Stereo System
To further assess the practicality and robustness of our DepthVanish patch, we evaluate its performance on a commercial stereo camera system, specifically the Intel RealSense D435i depth camera. We focus on evaluating the patch’s robustness from three aspects: model generalization, viewing orientation, and distance variation. ❶ Model generalization: we deploy the patches that optimized with five stereo models over KITTI dataset and evaluate their attack effective against D435i camera. As shown in Fig. 7, the patch consistently disrupts D435i predictions regardless of which model is optimized for, demonstrating strong attack transferability. ❷ Orientation robustness: we physically rotate the patch along the X and Y axes (see Fig. 3(a)). As visualized in Fig. 8, the patch (optimized with PSMNet on KITTI) remains effective under different viewing angles, confirming its robustness to rotation. ❸ Distance robustness: our method also shows robustness under varying distances. Corresponding visual results are provided in the supplementary material.